# Counting abelian extensions by Artin–Schreier conductor

### v1

@@ Abstract.
Let $G$ be a finite abelian $p$-group. We count $G$-extensions of global rational function fields $\mathbb{F}_{q}(T)$ of characteristic $p$ by the degree of what we call their Artin–Schreier conductor. The corresponding (ordinary) generating function turns out to be rational. This gives an exact answer to the counting problem, and seems to beg for a geometric interpretation.
@@ 1. Introduction
Let $G$ be a finite abelian group and let $K$ be a local or global field of characteristic $p$. There are several natural divisors one can associate to an (étale) $G$-extension $L$ of $K$, such as the discriminant divisor $\operatorname{disc}(L)$, or the conductor (in the sense of class field theory) $\operatorname{cond}(L)$. It is natural to ask for the number of $G$-extensions of $K$ such that the discriminant divisor (or the conductor divisor) has a particular degree $n$.
@@
$F^{\operatorname{asc}}_{K}(X):=\frac{1}{|G|}\sum_{L\textnormal{ étale $G$-extension of }K}X^{\deg(\operatorname{asc}(L))}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\deg(\operatorname{asc}(L))=n\}|\cdot X^{n}\in\mathbb{Z}\llbracket X\rrbracket$
@@
is the Hasse–Weil zeta function associated to $K$.
@@
In general, the rationality of the generating function fails when replacing the Artin–Schreier conductor by the ordinary conductor or by the discriminant. (See Section 5.)
It might also be worth pointing out that the asymptotics are quite a bit “simpler” when counting extensions of $\mathbb{F}_{q}(T)$ by Artin–Schreier conductor than when counting by ordinary conductor or discriminant. (Compare Corollary 4.5(b) to [Lag15, Theorem 1.2] and [Pot24, Theorem 5.30], respectively.)
Using inclusion–exclusion, one can show that the generating functions counting only those $G$-extensions of $K=\mathbb{F}_{q}(T)$ which are fields are still rational. (See Remark 3.3.)
Acknowledgements. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) — Project-ID 491392403 — TRR 358 (project A4). The author is grateful to Kiran Kedlaya, Jürgen Klüners, Nicolas Potthast, and Béranger Seguin for helpful discussions and comments on an earlier draft.
@@ 2. Definitions
For any étale $G$-extension $L$ of $K$ and any place $P$ of $K$, let $I^{t}_{P}(L|K)\subseteq G$ be the $t$-th ramification group in upper numbering. (Since $G$ is abelian, this group is independent of the choice of prime of $L$ above $P$.)
@@
We write $\operatorname{disc}(L)$ for the relative discriminant divisor of $L|K$. For any $\operatorname{inv}\in\{\operatorname{asc},\operatorname{cond},\operatorname{disc}\}$, we define
$F^{\operatorname{inv}}_{K}(X):=\frac{1}{|G|}\sum_{L\textnormal{ étale $G$-extension of }K}X^{\deg(\operatorname{inv}(L))}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\deg(\operatorname{inv}(L))=n\}|\cdot X^{n}.$
@@
$F^{\operatorname{asc}}_{K_{P}}(X)=\frac{1}{|G|}\sum_{L\textnormal{ étale $G$-extension of }K_{P}}X^{\operatorname{lstjmp}_{P}(X)}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\operatorname{lstjmp}_{P}(L)=n\}|\cdot X^{n}.$
@@ 3. Class field theory
    By class field theory, we have a bijection between étale $G$-extensions $L$ of $K_{P}$ and maps $\varphi:\mathcal{O}_{P}^{\times}\times\widehat{}\mathbb{Z}\rightarrow G$. There are exactly $|G|$ maps (continuous homomorphisms) ^$\widehat{}\mathbb{Z}\rightarrow G$, so each map $\varphi:U_{P}^{1}\rightarrow G$ has exactly $|G|$ extensions to $\mathcal{O}_{P}^{\times}\times\widehat{}\mathbb{Z}$. For $k\geq 0$, we have the following equivalences:
@@
$F^{\operatorname{asc}}_{K}(X)$ $=\sum_{L}X^{\deg(\operatorname{asc}(L))}=\sum_{L}\prod_{P}X^{\operatorname{lstjmp}_{P}(L)\deg(P)}$
@@
The étale $G$-extension $L$ corresponding to a map $\varphi:\operatorname{Gal}(K^{\mathrm{ab}}|K)\rightarrow G$ is a field if and only if $\varphi$ is surjective. Using inclusion–exclusion over the image of $\varphi$, one can therefore write the generating function $F^{\operatorname{inv}}_{K,\textnormal{field}}(X)$ counting only field $G$-extensions of $K$ as a linear combination of the generating functions counting étale $G^{\prime}$-extensions of $K$ for all subgroups $G^{\prime}\subseteq G$. In particular, if all the generating functions counting étale extensions are rational functions, then so is the generating function counting field extensions.
@@ 4. Artin–Schreier conductors
Let $t$ be a uniformizer of $K$, so that $K=\mathbb{F}_{Q_{P}}((t))$, and let $\omega_{1},\dots,\omega_{d}\in\mathbb{F}_{Q_{P}}\subset K$ form an $\mathbb{F}_{p}$-basis of $\mathbb{F}_{Q_{P}}$. For any $a=(a_{i})_{i}\in\Gamma_{P}^{1}=\prod_{i\geq 1:\ p\nmid i}\mathbb{Z}_{p}^{d}$ with $a_{i}=(a_{i1},\dots,a_{id})\in\mathbb{Z}_{p}^{d}$, we let
@@
Let $G=\prod_{e\geq 1}C_{p^{e}}^{m_{e}}$ and $c_{0},c_{1},\dots$ as in Theorem 1.1, let $K=\mathbb{F}_{q}(T)$ be a rational function field of characteristic $p$, and let $Z_{K}(X)$ be the Hasse–Weil zeta function associated to $K$.
@@
    $Z_{K}(X)=\sum_{D\textnormal{ divisor on }\mathbb{P}^{1}(\mathbb{F}_{q})}X^{\deg(D)}=\frac{1}{(1-X)(1-qX)}$
    of $\mathbb{P}^{1}(\mathbb{F}_{q})$ can be written as an Euler product
@@
From the partial fraction decomposition of the generating function $F^{\operatorname{asc}}_{K}(X)$, one can extract an exact formula for the number of étale $G$-extensions $L$ of $K$ with $\deg(\operatorname{asc}(L))=n$. (See for example the proof of [FS09, Theorem IV.9].)
@@
    |$|\{L\textnormal{ étale $G$-extension of }K:\deg(\operatorname{asc}(L))=n\}|\sim Cq^{a^{\prime}n}\qquad\textnormal{ for }n\rightarrow\infty.$
@@
(Plugging in $(X_{0},X_{1},\dots,X_{e-1})=(X,X^{p},\dots,X^{p^{e-1}})$, we recover the generating function $F^{\operatorname{asc}}_{K_{P}}(X)$ for the last jump $\operatorname{lstjmp}_{P}(\varphi)=\operatorname{jmp}_{P,0}(\varphi)$.) The corresponding global multivariate generating function $F^{\operatorname{jmp}}_{K}(X_{0},\dots,X_{e-1})=\prod_{P\in M_{K}}F^{\operatorname{jmp}}_{K_{P}}(X_{0},\dots,X_{e-1})$ is again a finite product of values of the Hasse–Weil zeta function, and in particular rational.

### v2

@@ Abstract.
Let $G$ be a finite abelian $p$-group. We count étale $G$-extensions of global rational function fields $\mathbb{F}_{q}(T)$ of characteristic $p$ by the degree of what we call their Artin–Schreier conductor. The corresponding (ordinary) generating function turns out to be rational. This gives an exact answer to the counting problem, and seems to beg for a geometric interpretation.
@@
Keywords. Asymptotics of Galois groups, global function fields, class field theory, meromorphic continuations
@@ 1. Introduction
Let $G$ be a finite abelian group and let $K$ be a local or global field of characteristic $p$. Let $K^{\operatorname{sep}}$ be the separable closure of $K$.
In this paper, the term $G$-extension will refer to what is often called an étale $G$-extension, a sub-$G$-extension, or a $G$-torsor. (See Section 2.1 for the definition.) Since $G$ is abelian, isomorphism classes of $G$-extensions are in bijection with continuous (not necessarily surjective) homomorphisms from the absolute Galois group of $K$ to $G$.
There are several natural divisors one can associate to an (étale) $G$-extension $L$ of $K$, such as the discriminant divisor $\operatorname{disc}(L)$, or the conductor (in the sense of class field theory) $\operatorname{cond}(L)$. It is natural to ask for the number of $G$-extensions of $K$ such that the discriminant divisor (or the conductor divisor) has a particular degree $n$.
@@
$F^{\operatorname{asc}}_{K}(X):=\frac{1}{|G|}\sum_{\begin{subarray}{c}L\textnormal{ (étale) $G$-extension of }K\\ \textnormal{up to isomorphism}\end{subarray}}X^{\deg(\operatorname{asc}(L))}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\deg(\operatorname{asc}(L))=n\}|\cdot X^{n}\in\mathbb{Z}\llbracket X\rrbracket$
@@
is the Hasse–Weil zeta function of $\mathbb{P}^{1}_{\mathbb{F}_{q}}$.
@@
Using inclusion–exclusion, one can show that the generating functions counting only those $G$-extensions of $K=\mathbb{F}_{q}(T)$ which are fields (or, equivalently, counting surjective continuous group homomorphisms from the absolute Galois group of $K$ to $G$) are still rational. (See Remark 3.3.)
@@
It might be worth pointing out that the asymptotics are quite a bit “simpler” when counting extensions of $\mathbb{F}_{q}(T)$ by Artin–Schreier conductor than when counting by ordinary conductor or discriminant. (Compare Corollary 4.5(b) to [Lag15, Theorem 1.2] and [Pot24, Theorem 5.30], respectively.)
The rationality of the generating function $F^{\operatorname{asc}}_{K}(X)$ is rather surprising: generating functions naturally appearing in arithmetic statistics are rarely known (or even expected) to even admit a meromorphic continuation to the entire complex plane. Over function fields $K=\mathbb{F}_{q}(T)$, the author is aware of the following cases:
  • The generating function counting $C_{3}$-extensions of $K=\mathbb{F}_{3}(T)$ by discriminant or by ordinary conductor does not have a meromorphic continuation to the entire complex plane. (See Section 5.)
  • If $G\neq 1$ is a finite abelian group and the characteristic of $K$ does not divide $|G|$ (unlike what we assume throughout the rest of this paper), one can show that the generating function counting $G$-extensions of $K$ by discriminant, by ordinary conductor, or by sum of ramified points only has a meromorphic continuation to the entire complex plane if $G=C_{2}$.
  • The Shintani zeta function counting $\operatorname{SL}_{2}(\mathbb{F}_{q}[T])$-orbits of binary cubic forms with coefficients in $\mathbb{F}_{q}[T]$ by discriminant is rational. A similar statement holds for adelic Shintani zeta functions, which essentially allow imposing a finite number of congruence conditions on the cubic forms. (See [Dat87].) However, despite Levi’s parameterization of cubic extensions by cubic forms, it seems doubtful whether the generating function counting cubic extensions of $\mathbb{F}_{q}(T)$ by discriminant would be rational (or even have a meromorphic continuation to the entire complex plane), essentially because one would need to impose infinitely many congruence conditions.
Acknowledgements. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) — Project-ID 491392403 — TRR 358 (project A4).
The author would like to thank Kiran Kedlaya, Jürgen Klüners, Nicolas Potthast, and Béranger Seguin for helpful discussions and comments on an earlier draft. Moreover, the author is grateful to the anonymous referees for their careful reading and helpful comments and corrections.
@@ 2. Definitions
2.1. $G$-extensions
Let $G$ be a finite group and let $K$ be a field.
[…]
Isomorphism classes of $G$-extensions of $K$ are in bijection with $G$-conjugacy classes of continuous homomorphisms from the absolute Galois group $\Gamma_{K}$ to $G$.
Any Galois extension $L/K$ together with an isomorphism $i:\operatorname{Gal}(L/K)\stackrel{{\scriptstyle\sim}}{{\to}}G$ is a $G$-extension, and the Galois extensions correspond to the surjective homomorphisms $\Gamma_{K}\to G$.
2.2. Invariants
@@
For any (étale) $G$-extension $L$ of $K$ and any place $P$ of $K$, let $I^{t}_{P}(L|K)\subseteq G$ be the $t$-th ramification group in upper numbering. (Since $G$ is abelian, this group is independent of the choice of prime of $L$ above $P$.)
@@
We write $\operatorname{disc}(L)$ for the relative discriminant divisor of $L|K$.
2.3. Generating functions
For any $\operatorname{inv}\in\{\operatorname{asc},\operatorname{cond},\operatorname{disc}\}$, we define
$F^{\operatorname{inv}}_{K}(X):=\frac{1}{|G|}\sum_{\begin{subarray}{c}L\textnormal{ (étale) $G$-extension of }K\\ \textnormal{up to isomorphism}\end{subarray}}X^{\deg(\operatorname{inv}(L))}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\deg(\operatorname{inv}(L))=n\}|\cdot X^{n}.$
@@ 2.3. Generating functions
$F^{\operatorname{asc}}_{K_{P}}(X)=\frac{1}{|G|}\sum_{\begin{subarray}{c}L\textnormal{ (étale) $G$-extension of }K_{P}\\ \textnormal{up to isomorphism}\end{subarray}}X^{\operatorname{lstjmp}_{P}(L)}=\frac{1}{|G|}\sum_{n\geq 0}|\{L:\operatorname{lstjmp}_{P}(L)=n\}|\cdot X^{n}.$
@@ 3. Class field theory
    By class field theory, we have a bijection between (étale) $G$-extensions $L$ of $K_{P}$ and maps $\varphi:\mathcal{O}_{P}^{\times}\times\widehat{}\mathbb{Z}\rightarrow G$. There are exactly $|G|$ maps (continuous homomorphisms) ^$\widehat{}\mathbb{Z}\rightarrow G$, so each map $\varphi:U_{P}^{1}\rightarrow G$ has exactly $|G|$ extensions to $\mathcal{O}_{P}^{\times}\times\widehat{}\mathbb{Z}$. For $k\geq 0$, we have the following equivalences:
@@
$F^{\operatorname{asc}}_{K}(X)$ $=\frac{1}{|G|}\sum_{L}X^{\deg(\operatorname{asc}(L))}=\frac{1}{|G|}\sum_{L}\prod_{P}X^{\operatorname{lstjmp}_{P}(L)\deg(P)}$
@@
The (étale) $G$-extension $L$ corresponding to a map $\varphi:\operatorname{Gal}(K^{\mathrm{ab}}|K)\rightarrow G$ is a field if and only if $\varphi$ is surjective. Using inclusion–exclusion over the image of $\varphi$, one can therefore write the generating function $F^{\operatorname{inv}}_{K,\textnormal{field}}(X)$ counting only field $G$-extensions of $K$ as a linear combination of the generating functions counting (étale) $G^{\prime}$-extensions of $K$ for all subgroups $G^{\prime}\subseteq G$. In particular, if all the generating functions counting (étale) extensions are rational functions, then so is the generating function counting field extensions.
@@ 4. Artin–Schreier conductors
Let $t$ be a uniformizer of $K_{P}$, so that $K_{P}=\mathbb{F}_{Q_{P}}((t))$, and let $\omega_{1},\dots,\omega_{d}\in\mathbb{F}_{Q_{P}}\subset K_{P}$ form an $\mathbb{F}_{p}$-basis of $\mathbb{F}_{Q_{P}}$. For any $a=(a_{i})_{i}\in\Gamma_{P}^{1}=\prod_{i\geq 1:\ p\nmid i}\mathbb{Z}_{p}^{d}$ with $a_{i}=(a_{i1},\dots,a_{id})\in\mathbb{Z}_{p}^{d}$, we let
@@
Let $G=\prod_{e\geq 1}C_{p^{e}}^{m_{e}}$ and $c_{0},c_{1},\dots$ as in Theorem 1.1, let $K=\mathbb{F}_{q}(T)$ be a rational function field of characteristic $p$, and let $Z_{K}(X)$ be the Hasse–Weil zeta function of $\mathbb{P}^{1}_{\mathbb{F}_{q}}$.
@@
    $Z_{K}(X)=\sum_{D\textnormal{ divisor on }\mathbb{P}^{1}_{\mathbb{F}_{q}}}X^{\deg(D)}=\frac{1}{(1-X)(1-qX)}$
    of $\mathbb{P}^{1}_{\mathbb{F}_{q}}$ can be written as an Euler product
@@
From the partial fraction decomposition of the generating function $F^{\operatorname{asc}}_{K}(X)$, one can extract an exact formula for the number of (étale) $G$-extensions $L$ of $K$ with $\deg(\operatorname{asc}(L))=n$. (See for example the proof of [FS09, Theorem IV.9].)
@@
    |$|\{L\textnormal{ (étale) $G$-extension of }K:\deg(\operatorname{asc}(L))=n\}|\sim Cq^{a^{\prime}n}\qquad\textnormal{ for }n\rightarrow\infty.$
@@
(Plugging in $(X_{0},X_{1},\dots,X_{e-1})=(X,X^{p},\dots,X^{p^{e-1}})$, we recover the generating function $F^{\operatorname{asc}}_{K_{P}}(X)$ for the last jump $\operatorname{lstjmp}_{P}(\varphi)=\operatorname{jmp}_{P,0}(\varphi)$.) The corresponding global multivariate generating function $F^{\operatorname{jmp}}_{K}(X_{0},\dots,X_{e-1})=\prod_{P\in M_{K}}F^{\operatorname{jmp}}_{K_{P}}(X_{0}^{\deg(P)},\dots,X_{e-1}^{\deg(P)})$ is again a finite product of values of the Hasse–Weil zeta function, and in particular rational.